# secReferences in the Online Appendix

### v1

@@ Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time. Using data from 2,500 campaigns, we find that NGOs are eight times more likely to launch campaigns on Annual General Meeting (AGM) dates. These campaigns increase media exposure and stakeholder scrutiny, raising the likelihood of consumer boycotts and related shareholder proposals at the following AGM of the targeted firm but have no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM outcomes, revealing the trade-offs they face to drive corporate change.
@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels. Through tactics of “private politics”—such as boycotts and shaming campaigns—NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017). Many studies examine how campaigns by both large and small NGOs impact and delegitimate targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009). For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007). Similar campaigns have targeted companies across sectors, including Nike, Shell, Monsanto, and BNP Paribas. However, we know relatively little about NGOs’ objectives and trade-offs and how they vary over their life cycle, shaping their campaign strategies toward misbehaving firms and, ultimately, their capacity to drive societal change.
To illustrate how objectives vary over time and in the life of an NGO, our analysis spans a wide range of organizations, from smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club.‘ However, we face several challenges, particularly due to the scarcity of data on NGOs. Donation data for non-political nonprofits is difficult to access because of donor privacy concerns, [^1] Data exist for donations to political parties and charities thanks to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023). This paper differs from this literature in that it does not focus on political donations or the resulting campaigns, but rather on the incentives that NGOs face when deciding to campaign against a firm. and NGOs have much more lenient transparency requirements than publicly listed corporations, complicating the tracking of their fund usage.
The absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot be easily reduced to profit maximization (Maxwell, 2012; Fioretti, 2022). The literature emphasizing extreme behaviors of non-profits, with some contribute to the greater good (e.g., Daubanes and Rochet, 2019) whereas others are captured by corporate interests (e.g., Bertrand et al., 2021). Therefore, this paper conducts a forensic investigation that infers NGOs’ objectives from their reactions to external incentives, focusing on the primary—and often only—observable data on NGOs: their campaigns, which must necessarily be public and observable to influence firms and their stakeholders.
@@
The starting point of our analysis is that NGO campaigns are strategically timed. Two key factors influence an NGO’s ability to capture stakeholders’ attention. First, media coverage amplifies NGO campaigns, making shaming and boycotting efforts more effective. Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from workers, consumers, shareholders, suppliers, and the public. We demonstrate that these two factors create a trade-off: NGOs must choose between seeking influence or enhancing their reputation.
@@
We find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM. The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008). This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, which might proxy for the NGO’s reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017).
@@
However, a late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020). Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times. In contrast, campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
@@
We argue that NGOs navigate this tradeoff differently depending on their accummulated reputation. Employing over ten years of campaign data against the largest US corporations, we first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly. Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late. We also find no evidence that NGO selects different target firms as they gain more campaigning experience.
@@
Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past. We interpret this as a reduced-form proxy for the reputation the NGO has built over time. However, there are diminishing returns to targeting AGM dates: once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes by campaigning earlier. Conversely, we find that accumulating experience at campaigning outside AGMs does not help NGOs at swaying shareholder proposal votes.
@@
Empirically, our approach leverages the exogenous timing of AGM meetings, which in the U.S. remains relatively stable within firms over time due to Securities and Exchange Commission (SEC) regulations designed to give shareholders adequate time to review proposals before voting them, while controlling for time-varying confounding factors at the firm, industry, or campaign topic levels. Therefore, while current events or news breaks may influence an NGO’s decision to launch a campaign, they are exogenous to the AGM calendar. We employ these external factors shifting NGO’s objectives to identify the trade-off between influence and visibility and to study how their objectives shift as they accumulate reputation or brand capital (Hasan et al., 2022).
@@
Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
@@
We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009). In this context, waiting for the target firm’s AGM may be seen as a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny and that is salient for shareholders.
@@
Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, allowing them to build their reputation. Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
@@
This paper also contributes to the growing body of literature on firm objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (Hawn and Ioannou, 2016) . [^3] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024). The literature emphasizes that shareholders can influence firms through either voicing concerns or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024)–especially through voting (Cuñat et al., 2016; Gine et al., 2017)–while consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010). However, other agents may play a pivotal role. Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019). This paper sheds new light on NGO as a coordination mechanism across stakeholders by providing new insights into the functioning of NGOs and how stakeholders respond to their activism.
Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the reputations of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024). In contrast, our study focuses on how NGOs can enhance their own reputation by adequately timing their campaigns against misbehaving corporations.
@@ 2 Data and Summary Statistics
Our dataset integrates three sources. First, we use data from Sigwatch, a European consultancy tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^4] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, MacDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs). For those firms, ISS data include detailed information on shareholder proposals for these firms, such as proposals withdrawn before the AGM.
@@
The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders. We present the key takeaways from each analysis at the end of each section. The empirical approach is discussed in detail next.
@@ 3 Empirical Strategy
Our results highlight NGO experience as a crucial factor in this trade-off. Before interpreting our findings, we present data in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past. Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets. They also have similar numbers of yearly news mentions and environmental and social scores, with comparable counts of environmental and social campaigns. Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels. This correlation analysis shows that NGOs’ campaign targets do not change as they gain experience, although they are more likely to focus on environmental issues rather than social ones in later years as climate change grew in importance.
@@ 4 Timing Campaigns for Media Exposure
Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise. We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
The vector $\mathbf{X}_{n,m,y}$ captures the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
@@
Media coverage of NGO campaigns is substantially greater in the month of the AGM.

### v2

@@ Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time. Using data from 2,500 campaigns, we find that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer boycotts and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels. Through tactics of “private politics,” such as boycotts and shaming campaigns, NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017). Many studies examine how campaigns by both large and small NGOs impact targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009). For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007). Similar campaigns have targeted companies across various sectors, including Nike, Shell, Monsanto, and BNP Paribas. Despite the significance of these campaigns in shaping NGO strategies and influencing corporate behavior, we know relatively little about NGOs’ objectives, trade-offs, and how these evolve over the course of an NGO’s life.
Researchers addressing these questions face several challenges, particularly due to the scarcity of data on NGOs. For instance, donation data for non-political nonprofits is difficult to access because of donor privacy concerns. [^1] Data on donations to political parties and charities are more readily available due to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023). This paper differs from this literature in that it does not focus on the connection between NGO and politics, but rather on the incentives that NGOs face when deciding to campaign against a firm. NGOs are also subject to much less stringent transparency requirements than other firms such as publicly listed corporations, which complicates the tracking of their fund usage. Moreover, the absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot easily be reduced to profit maximization (Maxwell, 2012; Fioretti, 2022). To obviate to this problem, the literature emphasizes extreme behaviors of NGOs, with some contributing to the greater good (e.g., Daubanes and Rochet, 2019), while others are captured by corporate interests (e.g., Bertrand et al., 2021, 2023).
@@
The starting point of our analysis is that NGO campaigns are strategically timed. Drawing on campaign data from a wide range of NGOs—spanning smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club‘—we identify two key factors that influence an NGO’s ability to capture stakeholders’ attention. First, media coverage amplifies NGO campaigns, making efforts like shaming and boycotting more effective. [^2] Consistent with this, Hawn (2021) finds that media coverage of a firm’s irresponsible behavior can significantly amplify negative public perceptions (Shea and Hawn, 2019), potentially resulting in the failure of planned merger. Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from stakeholders such as workers, consumers, shareholders, suppliers, and the public. We demonstrate that these two factors create a trade-off regarding the strategic decision of campaign timing. NGOs must choose between maximizing influence on firms and enhancing their reputation.
@@
Employing over ten years of campaign data against the largest US corporations, we find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM. The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008). This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, thereby accumulating reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017). Since our estimation routine compares the timing of the campaigns of NGOs campaigning against multiple firms in a year, we can account for any confounding factors varying within an NGO (and target firms) across years through fixed effects, such as their size.
@@
A late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020). Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times. Campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
@@
We argue that NGOs navigate this tradeoff differently depending on their reputation. We first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly. Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late. We also find no evidence that NGOs select different target firms as they gain more campaigning experience.
@@
Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past. We interpret this as a reduced-form proxy for the reputation built by the NGO for fighting over issues. As we document diminishing returns to targeting AGM dates, once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes through early campaigns. Conversely, we find that accumulating experience at campaigning outside AGM dates does not help NGOs at swaying shareholder proposal votes.
@@
Empirically, our approach leverages exogenous variation in the timing of AGM meetings, which in the U.S. we show to be relatively stable within firms over time due to regulations by the Securities and Exchange Commission (SEC) designed to give shareholders adequate time to review proposals before voting them. Across analyses, our approach also allows us to exclude time-varying confounding factors at the firm or industry, and campaign topic levels. In particular, while current events and news may influence an NGO’s decision to launch a campaign, they cannot influence the AGM calendar. As a result, objectives are identified by the tradeoff between targeting a firm’s AGM date to increase their exposure and impacting the firm’s decisions by campaigning early in support of the news or event.
@@
Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type (reputation)—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
@@
We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009). In this context, waiting for the target firm’s AGM is a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny at a time that is salient for shareholders and decision makers.
@@
Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, reflecting a better reputation. Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
@@
This paper also contributes to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; George et al., 2023; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (e.g., Hawn and Ioannou, 2016; Hartzmark and Sussman, 2019; Hasan et al., 2022). [^4] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016; Gine et al., 2017)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010). However, other agents may play a pivotal role. Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019).
Therefore, this short paper sheds new light on NGOs as coordinating various stakeholders such as shareholders and consumers, empirically highlighting a set of tradeoffs that arise from stakeholder involvement (Porter and Kramer, 2011; Kaplan, 2020). We provide new insights into the incentives NGOs face and how they can help align a firm’s action with the interests of its stakeholders (McGahan and Pongeluppe, 2023). In this direction, the results also suggest that other stakeholders can gain privately (i.e., reputation) by engaging with a firm, bringing empirical evidence to theories suggesting that stakeholders would not contribute resources to an organization without the potential to claim at least some of the value created (cf., Barney, 2018; McGahan, 2023).
Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the performances and prosocial outcomes of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024). Our study turns the table by focusing on the incentives that non-profit firms face when campaigning against for-profit firms.
@@ 2 Data and Summary Statistics
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^5] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs). For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
@@
The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders. We present the key takeaways from each analysis at the end of each section. Next, we discuss the empirical approach in detail.
@@ 3 Empirical Strategy
Our results highlight NGO experience as a crucial factor in this trade-off. Before interpreting our findings, we present results in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past. Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets. Targets also have similar numbers of yearly news mentions, environmental and social scores, and suffered a similar number of environmental and social campaigns. Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels. This correlation analysis shows that NGOs’ campaign targets do not change as NGOs gain experience, although they are more likely to focus on environmental issues rather than social ones in later years due to the increasing prominence of environmental challenges over social ones.
@@ 4 Timing Campaigns for Media Exposure
Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with the variable $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise. We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
The vector $\mathbf{X}_{n,m,y}$ contains the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
@@
Media coverage of NGO campaigns is substantially greater during the target’s AGM.